# Information theory for data-driven model reduction in physics and biology

### v3

@@ VII Relevant variable discovery in cyanobacterial populations
We now demonstrate how variational IB may be used as an aid for collective variable discovery in situations where physical intuition may not be a useful guide – collective behavior of biological organisms (Supplementary Movie 2). Here, we ask what the most predictive variables are for predicting the evolution of populations of cyanobacteria (Synechococcus elongatus). The dynamics of the colonies are driven by several factors: growth and division of individual bacteria, translational motion of groups of bacteria as they are pushed by their neighbors, as well as the circadian oscillations within each bacterium (Fig. 7a). These oscillations are controlled by three Kai proteins  [79] and depend in particular on the ratios of the copy number of these proteins which can be tuned experimentally [80].
@@
We were provided with videos of 10 cyanobacteria colonies that were grown under various conditions that impact their dynamics. However, as a test of our method, we were blinded to these conditions until we had performed our analysis. The videos are sequences of fluorescent images, taken once per hour, which show the clock state of each individual bacteria visualized with a fluorescent marker EYFP driven by the kaiBC promoter. Here, we focus on collective variables which are predictive of the state of the interior of the colony and not the growth in area of the colonies. We therefore crop the images to the interiors of each colony (SI Fig. S13). This allows us to isolate the motion of individual bacteria and fluorescence oscillations.
@@
Our input to the variational IB neural network are these cropped images augmented with a time-lagged image of the same region (Fig. 7b-c). The purpose of this time lag is to make the dynamics Markovian: the intensity field oscillating, the observation at a single time point does not allow to determine whether the intensity is increasing or decreasing. In principle more time-lags may be necessary, and their number could be estimated from the behavior of the mutual information (see SI Section 3) or the transition entropy [63] . As these quantities are difficult to measure from images with limited computational resources, we keep only one lag: the state is given by pairs $X_{t}=\{I(\mathbf{x},t),I(\mathbf{x},t+\tau)\}$, where $\tau$ is the lag duration. Here we take $\tau=3$ hr and a prediction horizon $\Delta t=8$ hr, but find that different $\Delta t$ or $\tau$ does not change our results (SI Fig. S13).
@@
Figure 7: Discovering slow collective variables in cyanobacteria populations (a) Fluorescent images of cyanobacteria colonies labeled with EYFP driven by the kaiBC promoter, allowing the visualization of Kai protein transcription. The colonies are imaged as they undergo cell growth and oscillations in Kai expression associated with the circadian rhythm. (b) The “state” used for variational IB is the time-lagged intensity field with lag time $\tau$ (see SI Fig. S13 for details). (c) Variational IB embeddings of time-lagged images into two dimensions (see Supplementary Movie 2). Every line corresponds to one colony’s evolution in time. Note the apparent oscillations of different radii. Three-dimensional embedding is shown in SI Fig. S15. One large-radius (blue) and one small-radius (black) trajectory are highlighted. (d) Mutual information between the future state and the encoding, given by $I(X_{t+\Delta t},H_{t})$, for varying dimension of the latent space. Each small point represents one training instance of the variational IB model, while the large point shows the maximum estimated value. $I_{\max}=I(X_{t},X_{t+\Delta t})$ is the mutual information estimated for the full dynamics. (e) Mean radius versus synchronization parameter for each colony (see SI Section 8). VIB identifies clusters of cells characterized by high (blue) or low (black) synchronization. This clustering corresponds to differing theophylline concentrations across experiments. Within each experimental movie (each of which contains 2-3 colonies) the radii are mostly constant. (f) Sample time series of weakly (black) and highly (blue) synchronized colonies. We apply a slight Gaussian blur to better visualize the bacteria boundaries.
@@
With variational IB we compress the state $X_{t}$ into a latent variable $h$ of variable dimension. We train the neural network on the entire dataset of all 10 colonies simultaneously. The dynamics in latent space undergo clear oscillations, indicating that the relevant variables encode primarily the intensity fluctuations rather than, for example, the spatial locations of the bacteria. Notably, the trajectories are essentially two-dimensional, even when the encoding space is higher dimensional. This is reflected in the information retained about the future state, $I(X_{t+\Delta t},H_{t})$: increasing the dimension of the embedding space beyond two only marginally increases $I(X_{t+\Delta t},H_{t})$; this tells us “when to stop” (Fig. 7d). We independently verify this by using principal component analysis to characterize the geometry of embedded trajectories, and find that even in higher dimensions the trajectories occupy a two dimensional subspace (SI Fig. S10). In the following, we thus restrict our focus to $\dim H=2$.
@@
We noticed that there were notable differences in the radius of latent space oscillations from colony to colony, two of which are highlighted in (Fig. 7c). To understand this difference, we examined the original microscopy time series corresponding to both large and small latent radius (Fig. 7e-f) and found that while the large-radius sample showed clear, nearly uniform oscillations in intensity, the small-radius samples appeared much more heterogeneous.
@@
To quantify this we consider each pixel to be an independent oscillator, akin to a spatial Kuramoto model [81, 82, 83], and compute a global synchronization order parameter $r(t)$ (see SI Section 7). For each colony we calculate the time-averaged synchronization $\langle r(t)\rangle_{t}$ and find that two clusters emerge corresponding to high and low synchronization (Fig. 7e). These clusters are precisely those representing trajectories of large and small latent radius, suggesting that variational IB learns to encode the synchronization of the colony in the latent variable radius. As a check, we perform IB on a simulated locally-coupled Kuramoto model as a system sharing many features of the experimental system. Here we also learn an encoding in which the latent radius corresponds to the synchronization order parameter (SI Fig. S11).
@@
In the SI we compare the performance of variational IB to several other model reduction methods and find that IB delivers more interpretable and well-behaved features. This is likely due to the fact that many standard methods for data-driven model reduction rely on assumptions about the dynamics which may not be appropriate in the case at hand, such as linearity. Even among deep learning methods free of such assumptions, such as time-lagged autoencoders, the variables learned by IB appear more interpretable. This increased interpretability is likely due to the compression term which effectively regularizes the latent space by encouraging the network to learn slow transfer operator eigenfunctions. While there are many specific variants of DMD [17, 84, 85, 86, 87] or autoencoders for dynamics [28, 29, 30, 88] that may outperform variational IB in some cases, we find that in this real-world example it yields the smoothest and most interpretable latent variables without a tailored pre-processing (SI Fig. S17).
@@
VIII Conclusion
We have related information-theoretical properties of dynamical systems to the spectrum of the transfer operator. We illustrate our findings on several simple and analytically tractable systems, and turn them into a practical tool using variational IB, which learns an encoding variable with a neural network. The latent variables of these networks can be interpreted as transfer operator eigenfunctions even though the network was not explicitly constructed to learn these: it optimizes a purely information-theoretic objective that contains no knowledge of a transfer operator or dynamics. This allows one to harness the power of neural networks to learn physically-relevant latent variables and, combined with methods such as SINDY, even complete reduced models with their governing equations. Biological systems are an ideal setting for such methods: despite their apparent complexity, they can often be captured by low-dimensional descriptions which are difficult to identify by physical considerations alone [3, 62, 89, 90]. We have shown that variational IB is a potentially powerful tool for these cases, and can discover slow variables even directly from image data without significant preprocessing. In the SI we showcase further applications, including to chaotic systems, where they may yield computationally efficient methods to calculate e.g. fractal dimensions.
IX Methods
Information processing inequality
Here we show how to obtain (2) from (1). The variables we are considering form the Markov chain
$X_{t+\Delta t}-X_{t}-h_{\text{enc}}(X_{t})-\mathcal{S}^{\Delta t}h_{\text{enc}}(X_{t})-h^{-1}_{\text{dec}}(\mathcal{S}^{\Delta t}h_{\text{enc}}(X_{t}))$
@@ IX Methods
$I(h^{-1}_{\text{dec}}(\mathcal{S}^{\Delta t}h_{\text{enc}}(X_{t})),X_{t+\Delta t})$ $\leq I(\mathcal{S}^{\Delta t}h_{\text{enc}}(X_{t}),X_{t+\Delta t})$
$\leq I(h_{\text{enc}}(X_{t}),X_{t+\Delta t}),$
@@
where the inequalities are maximized if $h^{-1}_{\text{dec}}$ and $\mathcal{S}^{\Delta t}$ are both bijective functions. Notably, $\mathcal{S}^{\Delta t}=\text{identity}$ is a solution. However, this should not be taken to mean that $H_{t}$ has no dynamics, as it will still have dynamics induced by the evolution of $X_{t}$.
The information bottleneck
The information bottleneck [38] is an example of a rate-distortion problem which seeks to find an optimal compression which minimizes some distortion measure with the original signal [58]. Concretely, we call $X$ the source signal, and let $H$ denote the compressed signal. In IB, rather than using an a priori unknown distortion function, one seeks to ensure that the compression retains information about an additional relevance variable $Y$. As noted in the main text, the IB optimization objective is given by the Lagrangian
$\mathcal{L}_{\text{IB}}[p(h|x)]$ $=I(X,H)-\beta I(Y,H),$ (12)
where in our case the source signal $X$ is the state of the system $X_{t}$ at time $t$, and the relevance variable is the state of the system $X_{t+\Delta t}$ at a future time $t+\Delta t$. The encoder which optimizes this objective can be solved for exactly and is given by [38]

### v4

@@ V.2 Cyanobacteria: latent variable discovery in experimental data
Variational IB may be used as an aid for collective variable discovery in situations where physical intuition may not be a useful guide – collective behavior of biological organisms (Supplementary Movie 2). Here, we ask what the most predictive variables are for predicting the evolution of populations of cyanobacteria (Synechococcus elongatus). The dynamics of the colonies are driven by several factors: growth and division of individual bacteria, translational motion of groups of bacteria as they are pushed by their neighbors, as well as the circadian oscillations within each bacterium (Fig. 7a). These oscillations are controlled by three Kai proteins  [104] and depend in particular on the ratios of the copy number of these proteins which can be tuned experimentally [105].
@@
We were provided with videos of ten cyanobacteria colonies that were grown under different conditions.The videos are sequences of fluorescent images, taken once per hour, which show the clock state of each individual bacteria visualized with a fluorescent marker EYFP driven by the kaiBC promoter. Here, we focus on collective variables which are predictive of the state of the interior of the colony and not the growth in area of the colonies. We therefore crop the images to the interiors of each colony (SI Fig. S14) to isolate bacterial motion and fluorescence oscillations inside each colony .
@@
Our input to the variational IB neural network are these cropped images augmented with a time-lagged image of the same region (Fig. 7b-c). The purpose of this time lag is to make the dynamics Markovian: because the intensity field oscillates, an observation at a single time point does not allow to determine whether the intensity is increasing or decreasing. In principle more time-lags may be necessary, and their number could be estimated from the behavior of the mutual information or the transition entropy (Ref. [88] and SI Section S3). As these quantities are difficult to measure from images with limited computational resources, we keep only one lag: the state is given by pairs $X_{t}=\{I(\mathbf{x},t),I(\mathbf{x},t+\tau)\}$, where $\tau$ is the lag duration. Here we take $\tau=3$ hr and a prediction horizon $\Delta t=8$ hr, but find that different $\Delta t$ or $\tau$ does not change our results (SI Fig. S14).
@@
Figure 7: Discovering slow collective variables in cyanobacteria populations (a) Fluorescent images of cyanobacteria colonies labeled with EYFP driven by the kaiBC promoter, allowing the visualization of Kai protein transcription. Colonies are imaged as they undergo cell growth and Kai expression oscillations associated with the circadian cycle. (b) The “state” used for variational IB is the time-lagged intensity field with lag time $\tau$ (see SI Fig. S13 for details). (c) Variational IB embeddings of time-lagged images into two dimensions (see Supplementary Movie 2). Every line corresponds to one colony’s evolution in time. Note the apparent oscillations of different radii. One large-radius (blue) and one small-radius (black) trajectory are highlighted. (d) Mutual information between the future state and the encoding, given by $I(X_{t+\Delta t},H_{t})$, for varying dimension of the latent space. Each small point represents one training instance of the variational IB model, while the large point shows the maximum estimated value. $I_{\max}=I(X_{t},X_{t+\Delta t})$ is the mutual information estimated for the full dynamics. (e) Mean radius versus synchronization parameter for each colony (SI Section S9). VIB identifies clusters of cells characterized by high (blue) or low (black) synchronization. (f) Sample time series of weakly (black) and highly (blue) synchronized colonies. We apply a Gaussian blur to better visualize the bacteria boundaries.
@@
With variational IB we compress the state $X_{t}$ into a latent variable $h$ of variable dimension. We train the neural network on the entire dataset of all ten colonies simultaneously. The dynamics in latent space undergo clear oscillations, indicating that the relevant variables encode primarily the intensity fluctuations rather than, for example, the spatial locations of the bacteria. Notably, the trajectories are essentially two-dimensional, even when the encoding space is higher dimensional. This is reflected in the information retained about the future state, $I(X_{t+\Delta t},H_{t})$: increasing the dimension of the embedding space beyond two only marginally increases $I(X_{t+\Delta t},H_{t})$; this tells us “when to stop” (Fig. 7d and SI Fig. S15). In the following, we thus restrict our focus to $\dim H=2$.
@@
We noticed that there were notable differences in the radius of latent space oscillations from colony to colony, two of which are highlighted in Figure 7c. To understand this difference, we examined the original microscopy time series corresponding to both large and small latent radius (Fig. 7e-f) and found that while the large-radius sample showed clear, nearly uniform oscillations in intensity, the small-radius samples appeared much more heterogeneous.
@@
To quantify this we consider each pixel to be an independent oscillator, akin to a spatial Kuramoto model [106, 107, 108], and compute a global synchronization order parameter $r(t)$ (SI Section S8). For each colony we calculate the time-averaged synchronization $\langle r(t)\rangle_{t}$ and find that two clusters emerge corresponding to high and low synchronization (Fig. 7e). These clusters are precisely those representing trajectories of large and small latent radius, suggesting that variational IB learns to encode the synchronization of the colony in the latent variable radius. As a check, we perform IB on a simulated locally-coupled Kuramoto model as a system sharing many features of the experimental system. Here we also learn an encoding in which the latent radius corresponds to the synchronization order parameter (SI Fig. S16).
@@
In Appendix E and Fig. 8 we compare the performance of variational IB to several other model reduction methods and find that IB delivers more interpretable and well-behaved features. This is likely due to the fact that many standard methods for data-driven model reduction rely on assumptions about the dynamics which may not be appropriate in the case at hand, such as linearity. Even among deep learning methods free of such assumptions, such as time-lagged autoencoders, the variables learned by IB appear more interpretable. This increased interpretability is likely due to the compression term which effectively regularizes the latent space by encouraging the network to learn slow transfer operator eigenfunctions. While there are many specific variants of DMD [109, 110, 111, 112, 113] or autoencoders for dynamics [49, 48, 114, 115] that may outperform variational IB in some cases, we find that in this real-world example it yields the smoothest and most interpretable latent variables without tailored preprocessing (Appendix E and Fig. 8).
@@
VI Conclusion
We have introduced an information-theoretic framework for dimensionality reduction in dynamical systems where relevant latent variables are identified based on how predictive they are about the future state of the system. In addition to formulating the core compression problem and characterizing its solution, our information-theoretic framework provides diagnostic tools to directly assess the goodness of the dimensionality reduction. In particular, we provide guidelines to decide when to stop increasing the size of the latent space given an objective, and to assess how (non-)Markovian the resulting latent dynamics are.
Our latent variable discovery procedure can be implemented exactly for low-dimensional discrete systems, or approximately with deep neural networks that can be deployed on continuous real-world data. Importantly, our mathematical characterization of the solution the optimization problem allows us to interpret the discovered latent variables in terms of spectral properties of the transfer operator. With cyanobacteria experiments as a case study, we have shown how variational IB can discover latent variables directly from experimental image data without significant preprocessing. We also showed how to combine these tools with existing equation-learning methods such as SINDY to produce a fully-fledged data-driven model discovery pipeline. In the SI we showcase further applications, including to chaotic systems, where they may yield computationally efficient methods to calculate chaotic invariants such as fractal dimensions directly from experimental data.
Going forward, we expect our framework to find uses in domains ranging from climate sciences where an overarching goal is to decouple as much as possible long-term climate from short-term weather despite the fact that many time scales are entangled [116, 34] to systems biology where, despite their apparent complexity, systems ranging from single cells or cell collectives [117, 118, 119, 120, 121, 122] to entire organisms like plants [123, 124] and animals [125, 86, 126] can often be captured by low-dimensional descriptions which are difficult to identify by biological or physical considerations alone.
Acknowledgments
Appendix A Mutual information as a measure of prediction error
Here we show that mutual information maximization emerges naturally as a measure of non-commutativity of the diagram Eq. (6) for probabilistic encoder/decoders. Let us measure the distance between the true distribution of future states and the decoded distribution of future states as
$d_{x,h^{\prime}}=D_{\text{KL}}\left[p(X_{t+\Delta t}|X_{t}=x)\Big\lVert p(X_{t+\Delta t}|H_{t+\Delta t}=h^{\prime})\right]$ (23)
This compares how well the latent future state $h^{\prime}$ predicts the distribution of $X_{t+\Delta t}$ compared to what would be obtained by evolving the true initial state $x$ forward in time. Note that $h^{\prime}$ follows the distribution $p(h^{\prime}|x)=\int\text{d}h_{t}\,p(h^{\prime}|h_{t})p(h_{t}|x)$ which results from first encoding $X_{t}$ to $H_{t}$ and then evolving $H_{t}$ to $H_{t+\Delta t}$.
The above Kullback-Leibler divergence compares the distributions for one particular pair of values $x$ and $h^{\prime}$. However, we are generally interested in the average distance over all possible values:
$d=\mathbb{E}_{p(X_{t},H_{t+\Delta t})}\left[d_{x,h^{\prime}}|X_{t}=x,H_{t+\Delta t}=h^{\prime}\right].$ (24)
This quantity can be written in terms of mutual information:
$d$ $=\mathbb{E}_{p(X_{t},H_{t+\Delta t},X_{t+\Delta t})}\left[\log\frac{p(X_{t+\Delta t}|X_{t})}{p(X_{t+\Delta t}|H_{t+\Delta t})}\right]$ (25)
$=\mathbb{E}_{p(X_{t},X_{t+\Delta t})}\left[\log\frac{p(X_{t+\Delta t}|X_{t})}{p(X_{t+\Delta t})}\right]-\mathbb{E}_{p(X_{t+\Delta t},H_{t+\Delta t})}\left[\log\frac{p(X_{t+\Delta t}|H_{t+\Delta t})}{p(X_{t+\Delta t})}\right]$ (26)
$=I(X_{t},X_{t+\Delta t})-I(H_{t+\Delta t},X_{t+\Delta t})$ (27)
To write the first line, we used the definition of the Kullback-Leibler divergence and the fact that variables follow the Markov chain $H_{t+\Delta t}-X_{t}-X_{t+\Delta t}$, which allowed us to write the expectation over the full joint distribution. To get the middle line we first multiplied the argument of the logarithm with $1=p(X_{t+\Delta t})/p(X_{t+\Delta t})$ and then integrated over $H_{t+\Delta t}$ in the first term and $X_{t}$ in the second. The final result shows that minimizing $d$ is equivalent to maximizing the mutual information $I(H_{t},X_{t+\Delta t})$ because the first term is fixed.
The above derivation does not incorporate the fact that we use a stochastic decoder $\tilde{h}^{-1}$ which follows the distribution $q(X_{t+\Delta t}|H_{t+\Delta t})$. Instead, we implicitly assumed that we could obtain the predicted distribution $p(X_{t+\Delta t}|H_{t+\Delta t})$ by performing the decoding $H_{t+\Delta t}\rightarrow H_{t}\rightarrow X_{t}\rightarrow X_{t+\Delta t}$ using Bayes’ rule. However, in some cases we may want to use a decoder which may be imperfect (for example because it is more computationally tractable). The distance $d$ resulting from using our imperfect decoder $q$ can be derived similarly to the above
$d$ $=\mathbb{E}_{p(X_{t},H_{t+\Delta t})}\left[D_{\text{KL}}\left[p(X_{t+\Delta t}|X_{t})\Big\lVert q(X_{t+\Delta t}|H_{t+\Delta t})\right]\right]$ (28)
$=\mathbb{E}_{p(X_{t},H_{t+\Delta t},X_{t+\Delta t})}\left[\log\frac{p(X_{t+\Delta t}|X_{t})}{q(X_{t+\Delta t}|H_{t+\Delta t})}\right]$ (29)
$=I(X_{t},X_{t+\Delta t})-I(H_{t+\Delta t},X_{t+\Delta t})+\mathbb{E}_{p(H_{t+\Delta t})}\left[D_{\text{KL}}\left[q(X_{t+\Delta t}|H_{t+\Delta t})\Big\lVert p(X_{t+\Delta t}|H_{t+\Delta t})\right]\right]$ (30)
$\equiv I(X_{t},X_{t+\Delta t})-I(H_{t+\Delta t},X_{t+\Delta t})+\delta_{\text{dec}}[h,\tilde{h}^{-1}]$ (31)
This distance is now given by the mutual information as before, plus a term $\delta_{\text{dec}}$ that measures the quality of our decoder. We retain its dependence on the (stochastic) encoder $h$ and decoder $\tilde{h}^{-1}$ for the next section. This term is greater than or equal to zero, and is minimized when the decoder step gives the same result as decoding $H_{t+\Delta t}\rightarrow H_{t}\rightarrow X_{t}\rightarrow X_{t+\Delta t}$ with Bayes’ rule. It is therefore known what decoder minimizes $\delta_{\text{dec}}$ (and hence which minimizes $d$). This is in contrast to the situation for the encoder $h$, whose optimal solution is unknown a priori.
Appendix B Separating concerns with the data processing inequality
Here we show how to obtain Eq. (10) in the main text. The starting point is Eq. (7), represented here explicitly
$h,\mathcal{V}_{\Delta t},\tilde{h}^{-1}$ $=\operatorname*{arg\,min}_{h,\mathcal{V}_{\Delta t},\tilde{h}^{-1}}\dim h(X_{t})+\beta\,d\left((\tilde{h}^{-1}\circ\mathcal{V}_{\Delta t}\circ h)(X_{t})),X_{t+\Delta t}\right)$ (32)
$=\operatorname*{arg\,min}_{h,\mathcal{V}_{\Delta t},\tilde{h}^{-1}}\mathcal{L}(h,\mathcal{V}_{\Delta t},\tilde{h}^{-1}).$ (33)
Here we write $h,\mathcal{V}_{\Delta t},\tilde{h}^{-1}$ and their composition as functions. These functions may be stochastic, which means that the their value is sampled from a distribution, e.g. $H_{t+\Delta t}=(\mathcal{V}_{\Delta t}\circ h)(X_{t})$ is drawn from the distribution
$p(h_{t+\Delta t}|x_{t})=\int\text{d}h_{t}\,\underbrace{p(h_{t+\Delta t}|h_{t})}_{\text{\ltxml@oqmark@open\textquotedblleft\penalty 10000\thinspace\thinspace$\mathcal{V}_{\Delta t}$\textquotedblright\ltxml@oqmark@close{}}}\underbrace{p(h_{t}|x_{t})}_{\text{\ltxml@oqmark@open\textquotedblleft\penalty 10000\thinspace\thinspace$h$\textquotedblright\ltxml@oqmark@close{}}}.$
There are two ways to introduce a distance function which compares the statistics of the decoded future state and the true future state. The first is simply to set $d=-I\left((\tilde{h}^{-1}\circ\mathcal{V}_{\Delta t}\circ h)(X_{t})),X_{t+\Delta t}\right)$. The second is to compare the distributions of decoded future states and true future states my measuring the Kullback-Leibler divergence between them, as we do in Appendix A. Both of these perspectives lead to the same result. For now, we focus on the first approach and comment on the second approach after.
By replacing minimization of the distance $d$ with maximization of the mutual information, we have
$\mathcal{L}_{\text{MI}}(h,\mathcal{V}_{\Delta t},\tilde{h}^{-1})$ $=\dim h(X_{t})-\beta\,I\left((\tilde{h}^{-1}\circ\mathcal{V}_{\Delta t}\circ h)(X_{t})),X_{t+\Delta t}\right).$ (34)
The variables in the above equation form the Markov chain
$X_{t+\Delta t}-X_{t}-h(X_{t})-\mathcal{V}_{\Delta t}(h(X_{t}))-\tilde{h}^{-1}(\mathcal{V}_{\Delta t}(h(X_{t})))$
@@ Appendix B Separating concerns with the data processing inequality
$I(\tilde{h}^{-1}(\mathcal{V}_{\Delta t}(h(X_{t}))),X_{t+\Delta t})\leq I(\mathcal{V}_{\Delta t}(h(X_{t})),X_{t+\Delta t})\leq I(h(X_{t}),X_{t+\Delta t}),$
@@
where the inequalities are maximized if $\tilde{h}^{-1}$ and $\mathcal{V}_{\Delta t}$ are both bijective functions. Notably, $\mathcal{V}_{\Delta t}=\text{identity}$ is a solution. However, this should not be taken to mean that $H_{t}$ has no dynamics, as it will still have dynamics induced by the evolution of $X_{t}$. Consequently,
$\min_{h,\mathcal{V}_{\Delta t},\tilde{h}^{-1}}\mathcal{L}_{\text{MI}}(h,\mathcal{V}_{\Delta t},\tilde{h}^{-1})$ $=\min_{h,\mathcal{V}_{\Delta t},\tilde{h}^{-1}}\dim h(X_{t})-\beta\,I\left((\tilde{h}^{-1}\circ\mathcal{V}_{\Delta t}\circ h)(X_{t})),X_{t+\Delta t}\right)$ (35)
$=\min_{h}\left(\dim h(X_{t})-\beta\,\min_{\mathcal{V}_{\Delta t},\tilde{h}^{-1}}I\left((\tilde{h}^{-1}\circ\mathcal{V}_{\Delta t}\circ h)(X_{t})),X_{t+\Delta t}\right)\right)$ (36)
$=\min_{h}\left(\dim h(X_{t})-\beta\,I\left(h(X_{t})),X_{t+\Delta t}\right)\right).$ (37)
This shows how the only quantity we need to learn is the encoder $h$. The latent evolution $\mathcal{V}_{\Delta t}$ and decoder $\tilde{h}^{-1}$ which minimize $\mathcal{L}_{\text{MI}}$ are induced by the encoding. While it may be difficult to numerically approximate these functions, they are at least known in principle, which is not the case for $h$.
We now illustrate how measuring the distance $d$ as a Kullback-Leibler divergence between the true distribution of future states and the predicted one leads to the same result. Plugging the distance $d$ derived in Appendix A into the full objective yields
$\mathcal{L}^{\prime}_{\text{MI}}(h,\mathcal{V}_{\Delta t},\tilde{h}^{-1})$ $=\dim h(X_{t})-\beta\,I\left((\mathcal{V}_{\Delta t}\circ h)(X_{t}),X_{t+\Delta t}\right)+\beta\delta_{\text{dec}}[h,\tilde{h}^{-1}].$ (38)
Because of the same Markov chain property used above, we have
$I(\mathcal{V}_{\Delta t}(h(X_{t})),X_{t+\Delta t})$ $\leq I(h(X_{t}),X_{t+\Delta t}),$
Consequently,
$\min_{h,\mathcal{V}_{\Delta t},\tilde{h}^{-1}}\mathcal{L}^{\prime}_{\text{MI}}(h,\mathcal{V}_{\Delta t},\tilde{h}^{-1})$ $=\min_{h,\mathcal{V}_{\Delta t},\tilde{h}^{-1}}\dim h(X_{t})-\beta\,I\left((\mathcal{V}_{\Delta t}\circ h)(X_{t}),X_{t+\Delta t}\right)+\beta\delta_{\text{dec}}[h,\tilde{h}^{-1}]$ (39)
$=\min_{h}\left(\dim h(X_{t})-\beta\,\min_{\mathcal{V}_{\Delta t}}I\left((\mathcal{V}_{\Delta t}\circ h)(X_{t}),X_{t+\Delta t}\right)+\beta\min_{\tilde{h}^{-1}}\delta_{\text{dec}}[h,\tilde{h}^{-1}]\right)$ (40)
$=\min_{h}\left(\dim h(X_{t})-\beta\,I\left(h(X_{t})),X_{t+\Delta t}\right)\right).$ (41)
To go from the second to last line to the last line, we used the data processing inequality to minimize the second term. For the last term $\delta_{\text{dec}}$, we used the fact that for any encoder $h$, the optimal decoder (which yields $\delta_{\text{dec}}=0$) is known (see Appendix A). In particular, it is
$p(x_{t+\Delta t}|h_{t+\Delta t})=\int\text{d}h_{t}\text{d}x_{t}p(x_{t+\Delta t}|x_{t})p(x_{t}|h_{t})p(h_{t}|h_{t+\Delta t})$ (42)
which is obtained by inverting the stochastic dynamics and the encoding using Bayes’ rule. Therefore, the only unknown quantity is $h$: all the others are induced by a choice of $h$.
Appendix C A primer on the information bottleneck
The information bottleneck was originally formulated in Ref. [41] as a rate distortion problem. Rate distortion theory describes how to maximally compress a signal (i.e. minimize its communication rate) such that it remains minimally distorted, which, however, presupposes the knowledge of a distortion function specifying which features need to be preserved [82]. In Ref. [41], Tishby et al. introduce a variant of the problem where the a priori unknown distortion function is not used, but rather one seeks to ensure that the compression retains information about an auxiliary variable $Y$ correlated with the signal. As the correlations with $Y$ define implicitly the relevant features of the signal to be preserved, $Y$ is called the relevance variable. Concretely, we call $X$ the source signal, and $H$ denotes the compressed signal. The random variables form a Markov chain HXYH-X-Y, meaning that $H$ and $Y$ are conditionally independent given $X$:
$p(y|h,x)$ $=p(y|x)p(h,x)$
$p(h|x,y)$ $=p(h|x)p(y,x)$
$\mathcal{L}_{\text{IB}}[p(h|x)]$ $=I(X,H)-\beta I(Y,H).$ (43)
$\begin{split}\mathcal{L}_{\text{IB}}[p(h|x)]&=\sum_{x,h}p(h|x)p(x)\log\frac{p(h|x)}{p(h)}-\beta\sum_{y,h}p(y,h)\log\frac{p(y,h)}{p(h)p(y)}+\sum_{x}\lambda(x)\left(1-\sum_{h}p(h|x)\right).\end{split}$ (44)
$\frac{\delta}{\delta p(h|x)}p(h^{\prime})$ $=\delta(h-h^{\prime})p(x)$
which can be rearranged to give the optimal encoder